Colson Corp. had $900,000 net income in 2018. On January 1, 2018 there were 600,000 shares of common stock outstanding. On March 1, 400,000 shares were issued and on September 1, Colson bought 100,000 shares of treasury stock. The tax rate is 40%. In addition, Colson issued $2,000,000 of 6% convertible bonds at face value during 2017. Each $1,000 bond is convertible into 100 shares of common stock. No bond was converted into common stock in 2018. Required: a) Compute basic earnings per share for 2018. b) Compute diluted earnings per share for 2018.
Answer #1

a. Weighted average outstanding common shares = 600000+400000*10/12 - 100000*4/12

= 900000

Basic earnings per share = Net income / weighted average outstanding common share

= 900000/900000

= $1 per share

b. Diluted earnings per share = (net income + interest net of tax)/(common shares + convertible shares)

= (900000 + 2000000*6%*60% )/(900000+2000000*100/1000)

= $0.88 per share